12) Refer to the table below. Marthas opportunity cost of making of a pie is: Time to Time to Make a Pie Make a Cake 60 minuWhy is the answer C) for both instead of A) if opportunity cost is what you give up over what you make? For question 1 Martha is giving up time making cake to make 1 pie so why would it not be 80/60=4/3 not 3/4? Please explain, thank you in advance.
Answer #1

ANSWER:

1) Martha opportunity cost of making a pie = 80 / 60 = 4 / 3 of a cake. option a is the right answer.

2) Martha opportunity cost of making a cake = 60 / 80 = 3 / 4 of a pie. option a is the right answer.

The answers for both the parts is a. (please check with your teacher.)